you now have the mass of chlorine that combine with 1.00g of Fe for each of the two different compounds. Do the ratio of 1:2 support or obey the law of multiple proportions.
1
Expert's answer
2020-09-29T06:12:50-0400

"n(Fe) = m(Fe)\/A_r(Fe) = 1.00g\/55.845g\/mol=0.018mol"

"n(Cl) = 2n(Fe) = 2*0.018 mol = 0.036mol"